OpenStudy (anonymous):

can someone explain to me how would i graph y < x or y greater than or equal to x

OpenStudy (anonymous):

y < x when graphing inequalities make sure to graph it as you would normally. ie y<x graph the equation y=x then shade the values that are above the line since you're looking for values that are greater than x

OpenStudy (anonymous):

greater than or equal to would mean points on the line (line iteself) is also included in the solution set

OpenStudy (anonymous):

\[y \ge x \]

OpenStudy (anonymous):

how would i graph that

OpenStudy (anonymous):

same way :) y=x but this time shading everything above y=x AND the line itself is included in the value.

OpenStudy (anonymous):

and also for absolute value, if it's y = |x+1| - 4 and i plug in the numbers for x (-2,-1,0,1,2) would i make them automatically positive since its absolute value or make it positive after I solve the whole equation

OpenStudy (anonymous):

y = |-2 +1| - 4 or make them positive y = |2+1| - 4

OpenStudy (anonymous):

right so an absolute value graph is shaped like a V because there should be no negative values. However like all graphs it can be shifted up or down left and right. Any value INSIDE the absolute value will emerge positive before any further evaluation. so in your question y = |-2 +1| - 4 y = |-1| - 4 y = 1 - 4 y = -3

OpenStudy (anonymous):

alright!

OpenStudy (anonymous):

yep! if you thought I was helpful fan me! I prioritize all my help towards fans.

OpenStudy (anonymous):

y<x y-x<0 ....(1) this line passes through origin,hence take values of x other than 0 put x=1,y=2 2-1<0 1<0 which is not true. Hence(1,2) does not lie on the graph. 

OpenStudy (anonymous):

correction values of x and y

OpenStudy (anonymous):

O SHOOT IM WRONG. I meant values under x SORRY.